# Cheap Learning: Maximising Performance of Language Models for Social Data Science Using Minimal Data

### v1

@@ Abstract
The field of machine learning has recently made significant progress in reducing the requirements for labelled training data when building new models. These ‘cheaper’ learning techniques hold significant potential for the social sciences, where development of large labelled training datasets is often a significant practical impediment to the use of machine learning for analytical tasks. In this article we review three ‘cheap’ techniques that have developed in recent years: weak supervision, transfer learning and prompt engineering. For the latter, we also review the particular case of zero-shot prompting of large language models. For each technique we provide a guide of how it works and demonstrate its application across six different realistic social science applications (two different tasks paired with three different dataset makeups). We show good performance for all techniques, and in particular we demonstrate how prompting of large language models can achieve high accuracy at very low cost. Our results are accompanied by a code repository to make it easy for others to duplicate our work and use it in their own research. Overall, our article is intended to stimulate further uptake of these techniques in the social sciences.
@@ 1 Introduction
The systematic analysis of large-scale text data is a growing part of social science research. The rapid digitisation of social life has meant that daily activity has increasingly left digital ‘traces’ (Boyd and Crawford, 2011). Across the social sciences, many of the core objects of disciplinary study can be interrogated through analysis of large datasets created by these traces. Examples abound: large-scale corpora of news articles; datasets of political manifestoes and speeches; textual interactions between individuals around romantic relationships, job searches, discussions around health and climate change, to name but a few (Scharkow and Vogelgesang, 2011; Burscher et al., 2015; Beigman Klebanov and Madnani, 2020; Gangula et al., 2019; Souma et al., 2019; Stede and Patz, 2021).
@@
In the past, large-scale text corpora were addressed through equally large-scale manual coding efforts (see, for example, Lehmann et al. (2023)), whereby texts are reviewed by teams of trained coders. However, around a decade ago, work emerged pointing to the potential for automatic content analysis to shed light on the social dynamics contained within these datasets, making use of insights from the field of machine learning (Wiedemann, 2013; Grimmer and Stewart, 2013; Grimmer et al., 2021). By now a considerable body of work has developed methods for automating content analysis in the social sciences (Nelson and Taneja, 2018; Németh et al., 2020), which offers obvious potential benefits in terms of saving time and effort, and thus allowing for analyses at a scale that would be beyond most teams of manual coders (Burscher et al., 2014).
@@
However, while the potential opened up by this automatic work is important, there remains a weakness in the automatic content analysis based on the ‘conventional’ paradigm of machine learning (as we will refer to it here): it still in reality contains a considerable manual component. As we will describe more fully below, in order for a classifier to be trained to identify a given concept, labelled training data is required, often in large quantities. Hence a significant time investment is required before such a classifier can be developed. Whilst such upfront work can be justified in industry contexts that may require millions of data points to be labelled automatically on an ongoing basis, many analytical research tasks in the social sciences could actually be achieved with this initially labelled dataset, rendering the training of the model somewhat irrelevant. Hence, the practical advantages of this type of machine learning approach in the social sciences is not always clear cut, with some saying that the time and financial savings from automated methods are exaggerated (De Grove et al., 2020; Németh et al., 2020) and others preferring methods granting researchers greater measurement validity (Baden et al., 2022).
@@
The heavy data requirements of conventional machine learning approaches have led to considerable research interest in techniques that might reduce data requirements without sacrificing performance or analytical rigour (Al-Jarrah et al., 2015; Goldsteen et al., 2022). In computer science, such techniques are now developing very actively, with considerable advances and a wide array of new techniques developing. However, only a few papers have emerged trying to make these techniques accessible to social science researchers (Molina and Garip, 2019; Macanovic, 2022; Wankmüller, 2022), and the uptake remains patchy at best.
@@
Each one of these techniques offers the potential for realising the benefits of supervised machine learning, but at less cost in terms of human labelling, hence offering the tantalising potential of finally realising the potential of automatic content analysis in the social sciences. For each one of these techniques, we describe how it functions, discuss approaches to validation, and discuss common potential use cases. In an empirical section, we then illustrate using existing annotated datasets how good performance can be achieved with a minimal amount of labelled data (up to 1024 data points), as well as showing how performance improvements relate to increases in labelled data. We also compare all of these techniques to the Naïve Bayes approach, a classic method from conventional machine learning Kibriya et al. (2005). Although the method was developed more than two decades ago, its simplicity of training and deployment makes it a good baseline to compare with newer and more sophisticated methods in a low-resource context.
@@
We should note the existence of a handful of other papers that have made similar contributions to our own (Terechshenko et al., 2020; Treviso et al., 2023). Whilst these are valuable works, our paper is distinctive in a number of ways. First, these papers address one technique exclusively, whereas ours compares three at the same time. This comparison allows us to explore the ‘labelling budget’ (the amount of manually labelled data required for each technique), as well as discussing how different methods perform well in different circumstances. Second, alongside our paper we release code that will allow other researchers with a practical working knowledge of the Python language to implement these techniques in their own research. [^1] https://github.com/Turing-Online-Safety-Codebase/cheap_learning
@@ 2.1 Weak Supervision
Labelling functions are abstract rules that take an unlabelled data point and either assign a label if a criterion is met or abstain from assigning it otherwise. The presence or absence of keywords contained in text data is one of the most obvious types of such rules, however, there is no need to only make use of keywords. Regular expressions, Named-Entity Recognition (NER) systems, results from other models such as sentiment classifiers, the length of the document to be classified, or its source; any of these document features can be incorporated into a labelling function. This diversity of potential criteria makes it possible to have topic-specific sets of labelling functions obtained from the particularities of the topic to be classified, going a step beyond the dictionary approach. Some examples of simple labelling functions that we make use of in the experimental section of the paper are presented below. For further examples, the list of all the labelling functions used for the classification tasks presented in the paper can be found in Appendix A.3.
@@
The exploration subset of the dataset serves to fine-tune or debug the set of labelling functions. Weak supervision enables measuring coverage and overlap of these functions in the different sets (details of the LFs used in this work are presented in Appendix A). Coverage for a specific labelling function indicates the percentage of data points it labels without abstaining. On the other hand, overlap represents the percentage of data points where two or more labelling functions do not abstain. Ideally, for a given class, we aim for a set of labelling functions that collectively cover the entire exploration set without significant redundancy in their actions, measured by a high overlap.
@@ 3.1 Data
Wikipedia Personal Attacks
All
402 734 10000 1
Positive
341 938 10000 5
Negative
410 702 10000 1
IMDb Movie Review Sentiment
All
1297 980 14106 32
Positive
1303 1014 13593 65
Negative
1271 928 8698 32
@@
Dataset Split Label All Positive Negative Wikipedia Personal Attacks All 115,864 13,590 102,274 Train 69,526 8,079 61,447 Test 23,178 2,756 20,422 Development 23,160 2,755 20,405 Exploration (WS) 100 12 88 IMDb Movie Review Sentiment All 50,000 25,000 25,000 Train 25,000 12,500 12,500 Test 12,500 6,250 6,250 Development 12,500 6,250 6,250 Exploration (WS) 100 92 8
@@ 3.2 Model Selection and Training Details
For the conventional baseline models, we develop a Multinomial Naive Bayes classifier (typically used for text classification - e.g., to represent how many times a word appears in a document) built based on Bayes’s theorem (Kibriya et al., 2005). [^5] We use the package developed by scikit-learn: https://scikit-learn.org/stable/modules/generated/sklearn.naive_bayes.MultinomialNB.html This method is selected because it can be easily developed and speedily trained without using GPUs. We use TF-IDF to convert data into vectors/features.
@@ 4 Results
In this section, we present the results of our experiments, to illustrate these techniques in action. Although our experiments made use of multiple models and configurations, in order to simplify presentation we display just a sample of these. In particular, for transfer learning we show results generated using DistilBERT, and for prompt engineering we also look at DistilBERT and look at only one fixed prompt (‘Does this text contain abuse?’ for abuse classification and ‘Does this text contain negative sentiment?’ for movie sentiment classification). For Naïve Bayes and weak supervision, we present the full results. Other base models and prompts provided similar results.
We begin by investigating our binary abuse classification task, shown in Figure 1. This figure has several panels, reflecting whether the underlying data is balanced or not. We can see that in the simple case of underlying balanced data (Figure 1, Panel (a)), which will hence generate both balanced training datasets and test datasets), there is actually little difference between any of the techniques in question. The ‘conventional’ machine learning approach achieves a high performance (Macro F1 of around 0.8) with only around 100 examples. While it is eventually outperformed by prompt engineering, transfer learning and weak supervision, this is only after the labelling budget crosses the 250 mark, and even after 1,000 examples have been used the difference is not enormous. Zero-shot learning with GPT-3.5 and GPT-4 offers an impressive Macro F1 (around 0.9) without any training data (when interpreting the labelling budgets, we should also bear in mind that, in addition to the training budget, we need to consider the 100 data points used in the exploration data set for the weak supervision labelling functions).
@@
However, we know that the prevalence of abusive text is generally lower than the presence of non-abusive text. Hence a more ‘realistic’ scenario is that both training and test data are unbalanced (Figure 1, Panel (b)). Here we can see that conventional machine learning performs much worse than its ‘cheaper’ counterparts, with a Macro F1 score of less than 0.5 for all levels of the training budget investigated. In the unbalanced case, where abuse is around 10% of the total dataset, then even with a relatively large labelling budget only a small amount of cases from the positive class will be seen. The difference between the other techniques in particular is notable, with transfer learning able to achieve the highest overall performance (though the Macro F1 of 0.8 from GPT-4 is again notable, considering that no training data is required to achieve it).
@@
Of course, in the case of unbalanced data, one common workaround is to artificially balance the training dataset (though of course preserving the real underlying distributions of data in the test data). This can improve model performance on a limited training budget (though achieving training balanced data still requires a higher labelling effort, or trying to artificially inflate the likelihood of labelling positive data through the use of keywords). Results from this approach are presented in Figure 1, Panel (c). We can see that model performance is improved in all cases, reaching the range of 0.7-0.8 after a few hundred examples (the zero-shot approaches all fall within this range as well). Weak supervision is actually the best-performing technique in this scenario (as the test data is the same, the results for zero-shot prompt engineering are the same in Panels (b) and (c)).
@@
We will now move on to address our binary movie sentiment classification task, shown in Figure 2. The format of this figure is the same, with panels addressing classifier performance over a variety of labelling budgets for balanced data, unbalanced data, and a balanced training set with unbalanced test data. Again, these three setups were chosen as representative of common real-world research scenarios. The overall results from this task are similar to the abuse classification task. For balanced data, we can see that all of the methods perform in a similar fashion. Our conventional machine learning approach achieves a Macro F1 of 0.8 with around 100 labelled examples. Prompt engineering, transfer learning and weak supervision are able to outperform it, especially as the labelling budget increases, eventually arriving at scores of around 0.9. Again, zero-shot learning performs well, with performance above 0.9 for GPT-4. In the unbalanced case, the performance of all models is lower, though transfer learning and prompt engineering can achieve a Macro F1 of 0.8 with around 1,000 examples, again common with GPT-4. Our Naïve Bayes model performs very poorly in this context. Finally, when training data is balanced but test data remains unbalanced, more of the approaches can achieve a Macro F1 of around 0.8, though we would recall again that the cost of generating a balanced training dataset from unbalanced data can be quite high.
@@
The IMDB movie sentiment database (Maas et al., 2011) has been, as mentioned in the Methods section 3.1, extensively studied in the literature. As such, a natural worry arises that the database has been used in the training datasets of GPT-3.5 and GPT-4, thus contaminating the zero-shot testing described above. Given that we are unable to check if indeed the database was used to train the GPT models, we collected an analogous validation of movie reviews from another review aggregator, TMDB [^6] https://www.themoviedb.org/. The included reviews were publicly published from October 2021, past the training date cut-off for the GPT-3.5. As in Maas et al. (2011), we label as positive those reviews with a score higher than 6 out of 10, and we label negative those reviews leaving a score lower than 5, omitting neutral reviews and including no more than 30 reviews from a single movie. Our final database comprises 855 reviews, with a distribution of 73.3% of positive reviews and 26.7% of negative reviews. When performing a zero-shot test on the TMDB dataset, we obtain Macro F1 scores between 0.87 and 0.94 for both GPT-3.5 and GPT-4, using the three prompts as the other tasks. GPT-3 has, on the other hand, different performances depending on the prompt, with Macro F1 scores of 0.67, 0.85 and 0.11. The results are in accordance with the results presented in Figure 2, thus validating them for our use of the three versions of GPT. In other words, we are confident the high performance of GPT in these classification tasks is not solely a result of GPT’s training data already including the IMDB dataset.
@@
It is also worth evaluating the amount of training time required for each of the techniques, something which has knock-on consequences for the financial costs involved if making use of a cloud platform. Overall, the training time is short for all three techniques and two datasets (see Figures 3 and 4). As the labelling budget grows, the training time also grows linearly except for weak supervision. For weak supervision, increasing labelling examples does not seem to affect training time when trained on Wikipedia personal attacks dataset. The training time is generally around 1 minute regardless of labelling budgets and slightly decreases when using the maximum labelling budget in our experiments (n=1024). Training times in this case are a sum of two elements: the time that the set of labelling functions takes to label data points and the time that the probabilistic LabelModel (Ratner et al., 2019) takes to compute the weights and probabilities associated to the labelling functions. The different behaviours seen in Figure 3 and Figure 4 depend on the nature of the labelling functions used. In the case of the binary movie sentiment analysis, we add two regex LFs, thus observing the linear increase of time as they require more time to look for matches as the number of data points increases. In the case of the binary abuse experiment, the labelling functions used are dictionary-based or transforming annotators’ knowledge into a labelling function. Given that directory-based functions (looking for words in a text) are not as computationally expensive as regular expression functions, given that the maximum length of the text is fixed, we obtain a stable training time of 1 minute.
@@
While Zero-shot Prompt Engineering does not require any training time, it is worth briefly commenting on the costs that could be incurred through the API to perform classification tasks (these are correct at the time of writing). Costs through such an API are per ‘token’: a small piece of text that might be equivalent to a word, a part of a word or a piece of punctuation (as a rule of thumb, a token is often around four characters long). To perform a classification using an LLM there will be costs for both input tokens (the chunk of text to be classified) and output tokens (the response indicating the classification). Of course, the response is usually just one or two tokens long, meaning the length of the input text is the main factor determining API costs. At the time of writing, GPT-3 had a cost of £0.0006 per 1,000 tokens, GPT-3.5 a cost of £0.0013 per 1,000 tokens and GPT-4 a cost of £0.0025 per 1,000 tokens. The price for a project will, obviously, vary according to the model chosen and the type of input text but, to give an example, the movie reviews were on average around 325 tokens each, meaning that 10,000 movie reviews could be classified on GPT-4 for around £8. Hence, API costs are only likely to be a significant concern for relatively large pieces of text where there are millions of observations to be classified.

### v2

@@ Abstract
The field of machine learning has recently made significant progress in reducing the requirements for labelled training data when building new models. These ‘cheaper’ learning techniques hold significant potential for the social sciences, where development of large labelled training datasets is often a significant practical impediment. In this article we review three ‘cheap’ techniques that have developed in recent years: weak supervision, transfer learning and prompt engineering. For the latter, we also review the particular case of zero-shot prompting of large language models. For each technique, we provide a guide of how it works and demonstrate its application and the presence of systematic biases across two different and realistic social science tasks paired with three different dataset makeups. We show good performance for all techniques and we demonstrate how prompting of large language models can achieve high accuracy at very low cost, but biases must be considered.
@@ 1 Introduction
The systematic analysis of large-scale text data is a growing part of sociology and social science research. The rapid digitisation of social life has meant that daily activity has increasingly left digital ‘traces’ (Boyd and Crawford, 2011). Across the social sciences, many of the core objects of disciplinary study can be interrogated through analysis of large datasets created by these traces. Examples abound: large-scale corpora of news articles; datasets of political manifestoes and speeches; textual interactions between individuals around romantic relationships, job searches, discussions around health and climate change, to name but a few (Scharkow and Vogelgesang, 2011; Burscher et al., 2015; Beigman Klebanov and Madnani, 2020; Gangula et al., 2019; Souma et al., 2019; Stede and Patz, 2021; Jensen et al., 2022).
@@
In the past, large-scale text corpora were addressed through equally large-scale manual coding efforts (see, for example, Lehmann et al. (2023)), whereby texts are reviewed by teams of trained coders. However, around a decade ago, work emerged pointing to the potential for automatic content analysis to shed light on the social dynamics contained within these datasets, making use of insights from the field of machine learning (Wiedemann, 2013; Grimmer and Stewart, 2013; Grimmer et al., 2021). By now a considerable body of work has developed methods for automating content analysis in the social sciences (Nelson and Taneja, 2018; Németh et al., 2020; Lichtenstein and Rucks-Ahidiana, 2021), which offers obvious potential benefits in terms of saving time and effort, and thus allowing for analyses at a scale that would be beyond most teams of manual coders (Burscher et al., 2014).
@@
However, while the potential opened up by this automatic work is important, there remains a weakness in the automatic content analysis based on the ‘conventional’ paradigm of machine learning (as we will refer to it here): it still in reality contains a considerable manual component. As we will describe more fully below, in order for a classifier to be trained to identify a given concept, labelled training data is required, often in large quantities. Hence a significant time investment is required before such a classifier can be developed. Whilst such upfront work can be justified in industry contexts that may require millions of data points to be labelled automatically on an ongoing basis, many analytical research tasks in the social sciences could actually be achieved with this initially labelled dataset, rendering the training of the model somewhat irrelevant. Hence, the practical advantages of this type of machine learning approach in the social sciences are not always clear cut, with some saying that the time and financial savings from automated methods are exaggerated (De Grove et al., 2020; Németh et al., 2020) and others preferring methods granting researchers greater measurement validity (Baden et al., 2022).
@@
The heavy data requirements of conventional machine learning approaches have led to considerable research interest in techniques that might reduce data requirements without sacrificing performance or analytical rigour (Al-Jarrah et al., 2015; Goldsteen et al., 2022; Galal et al., 2024). In computer science, such techniques are now developing very actively, with considerable advances and a wide array of new techniques developing. However, only a few papers have emerged trying to make these techniques accessible to social science researchers (Molina and Garip, 2019; Macanovic, 2022; Wankmüller, 2022; Costantini et al., 2023), and the uptake remains patchy at best.
@@
Each one of these techniques offers the potential for realising the benefits of supervised machine learning, but at less cost in terms of human labelling, hence offering the tantalising potential of finally realising the potential of automatic content analysis in sociology and the social sciences in general. For each one of these techniques, we describe how it functions, discuss approaches to validation, and discuss common potential use cases. In an empirical section, we then illustrate using existing annotated datasets how good performance can be achieved with a minimal amount of labelled data (up to 1024 data points), as well as showing how performance improvements relate to increases in labelled data. We also compare all of these techniques to the Naïve Bayes approach and to a logistic regression classifier, two classic methods from conventional machine learning Kibriya et al. (2005); Aborisade and Anwar (2018). Although these methods were developed more than two decades ago, their simplicity of training and deployment makes them a good baseline to compare with newer and more sophisticated methods in a low-resource context.
We complement our performance analysis with an additional experiment to study the biases each technique can have in classifying text for each task. While the most recent OpenAI’s GPT models have been used to annotate text (Gilardi et al., 2023), the potential systemic biases they have at an annotation task can lead to incorrect results (Ashwin et al., 2023; Spirling, 2023). With this bias analysis, closed models like OpenAI’s GPTs can be compared to more classical machine learning techniques in another useful dimension for social scientists besides their performances.
@@
We should note the existence of a handful of other papers that have made similar contributions to our own (Terechshenko et al., 2020; Treviso et al., 2023). Whilst these are valuable works, our paper is distinctive in a number of ways. First, these papers address one technique exclusively, whereas ours compares three at the same time. This comparison allows us to explore the ‘labelling budget’ (the amount of manually labelled data required for each technique), as well as discussing how different methods perform well in different circumstances. Second, our analysis is not limited to the performance of labelling data but also studies the systematic bias of each technique. Finally, alongside our paper we release code that will allow other researchers with a practical working knowledge of the Python language to implement these techniques in their own research [^1] https://github.com/Turing-Online-Safety-Codebase/cheap_learning.
With this paper, we contribute by giving sociologists and social scientists tools and strategies to address challenges where data scarcity is found. By offering available online resources and a performance and bias analysis done on two well-known examples in the literature (Maas et al., 2011; Wulczyn et al., 2017), we hope for social scientists to apply these strategies in key challenges where data scarcity and bias might be embedded into the problem, like interview analysis of under-represented groups (Deterding and Waters, 2021; Small and Calarco, 2022; Bonikowski and Nelson, 2022; Ashwin et al., 2023; Atari et al., 2023) or text analysis and labelling using historical text archives (Jaillant, 2022; Marciano et al., 2018; Rahal et al., 2024).
While acknowledging the importance of unsupervised machine learning approaches (Hutto and Gilbert, 2014; Shelke et al., 2022), we chose not to work with these techniques. Given that unsupervised learning discovers structure in data rather than labelling according to prior theoretical concepts of interest, we consider unsupervised learning techniques working fundamentally different to the features of supervised learning we are interested in testing, thus falling out of the scope of this study. In the same way, we are also not including open-source large language models (LLMs) for this study. On the one hand, we are interested in showing the monetary costs, performance and potential biases of using the API of closed models like the OpenAI’s GPT-3.5 (ChatGPT) and GPT-4.0, which have grabbed much attention (Kalla et al., 2023) in the academic and non-academic world. On the other hand, we decided to centre our attention on the mentioned closed LLMs to avoid this study becoming another comparison between the different LLMs now available for researchers (Yu et al., 2023).
@@ 2.1 Weak Supervision
Labelling functions are abstract rules that take an unlabelled data point and either assign a label if a criterion is met or abstain from assigning it otherwise. The presence or absence of keywords contained in text data is one of the most obvious types of such rules, however, there is no need to only make use of keywords. Regular expressions, Named-Entity Recognition (NER) systems, results from other models such as sentiment classifiers, the length of the document to be classified, or its source; any of these document features can be incorporated into a labelling function. This diversity of potential criteria makes it possible to have topic-specific sets of labelling functions obtained from the particularities of the topic to be classified, going a step beyond the dictionary approach. Some examples of simple labelling functions that we make use of in the experimental section of the paper are presented below. For further examples, the list of all the labelling functions used for the classification tasks presented in the paper can be found in the Appendix A.4 of the online supplement.
@@
The exploration subset of the dataset serves to fine-tune or debug the set of labelling functions. Weak supervision enables measuring coverage and overlap of these functions in the different sets (details of the LFs used in this work are presented in Appendix A of the online supplement). Coverage for a specific labelling function indicates the percentage of data points it labels without abstaining. On the other hand, overlap represents the percentage of data points where two or more labelling functions do not abstain. Ideally, for a given class, we aim for a set of labelling functions that collectively cover the entire exploration set without significant redundancy in their actions, measured by a high overlap.
@@ 3.1 Data
Wikipedia Personal Attacks All 402 734 10000 1
Positive 341 938 10000 5
Negative 410 702 10000 1
IMDb Movie Review Sentiment All 1297 980 14106 32
Positive 1303 1014 13593 65
Negative 1271 928 8698 32
@@
Dataset Split Label All Positive Negative Wikipedia Personal Attacks All 115,864 13,590 (11.73%) 102,274 (88.27%) Train 69,526 8,079 (11.62%) 61,447 (88.38%) Test 23,178 2,756 (11.89%) 20,422 (88.11%) Development 23,160 2,755 (11.90%) 20,405 (88.10%) Exploration (WS) 100 12 (12%) 88 (88%) IMDb Movie Review Sentiment All 50,000 25,000 (50%) 25,000 (50%) Train 25,000 12,500 (50%) 12,500 (50%) Test 12,500 6,250 (50%) 6,250 (50%) Development 12,500 6,250 (50%) 6,250 (50%) Exploration (WS) 100 92 (92%) 8 (8%)
@@ 3.2 Model Selection and Training Details
For the conventional baseline models, we develop a Multinomial Naïve Bayes classifier (typically used for text classification - e.g., to represent how many times a word appears in a document) built based on Bayes’s theorem (Kibriya et al., 2005). [^5] We use the package developed by scikit-learn: https://scikit-learn.org/stable/modules/generated/sklearn.naive_bayes.MultinomialNB.html. A deeper analysis on why we choose the Multinomial Naïve Bayes and not other options like the presented in Rennie et al. (2003) and Fatourechi et al. (2008) can be read in the Appendix A.3 of the online supplement. For the logistic regression classifier, we construct a logistic regression with an L2 penalty parameter cross-validated to maximise the used F1 macro score [^6] We use the package developed by scikit-learn: https://scikit-learn.org/stable/modules/generated/sklearn.linear_model.LogisticRegression.html. These methods are selected because they can be easily developed and speedily trained without using GPUs. We use TF-IDF to convert data into vectors/features.
@@ 4 Results
In this section, we present the results of our experiments, to illustrate these techniques in action. Although our experiments made use of multiple models and configurations, in order to simplify presentation we display just a sample of these. In particular, for transfer learning we show results generated using DistilBERT, and for prompt engineering we also look at DistilBERT and look at only one fixed prompt (‘Does this text contain abuse?’ for abuse classification and ‘Does this text contain negative sentiment?’ for movie sentiment classification). For Naïve Bayes, logistic regression and weak supervision, we present the full results. Other base models and prompts provided similar results.
We begin by investigating our binary abuse classification task, shown in Figure 1. This figure has several panels, reflecting whether the underlying data is balanced or not. We can see that in the simple case of underlying balanced data (Figure 1, Panel (a)), which will hence generate both balanced training datasets and test datasets), there is actually little difference between any of the techniques in question. The ‘conventional’ machine learning approaches achieve a high performance (Macro F1 of around 0.8) with only around 150 examples. While it is eventually outperformed by prompt engineering, transfer learning and weak supervision, this is only after the labelling budget crosses the 250 mark777when interpreting the labelling budgets, we should also bear in mind that, in addition to the training budget, we need to consider the 100 data points used in the exploration data set for the weak supervision labelling functions, and even after 1,000 examples have been used the difference is not enormous. Zero-shot learning with GPT-3.5 and GPT-4 offers an impressive Macro F1 (around 0.9) without any training data.
@@
However, we know that the prevalence of abusive text is generally lower than the presence of non-abusive text. Hence a more ‘realistic’ scenario is that both training and test data are unbalanced (Figure 1, Panel (b)). Here we can see that conventional machine learning performs worse than its ‘cheaper’ counterparts, with a Macro F1 score of less than 0.5 for all levels of the training budget investigated for the Naïve Bayes classifier, and with high standard deviation and a lower Macro F1 score than the rest for the case of the logistic regression. In the unbalanced case, where abuse is around 10% of the total dataset, then even with a relatively large labelling budget only a small amount of cases from the positive class will be seen. The difference between the other techniques in particular is notable, with transfer learning able to achieve the highest overall performance (though the Macro F1 of 0.8 from GPT-4 is again notable, considering that no training data is required to achieve it).
@@
Of course, in the case of unbalanced data, one common workaround is to artificially balance the training dataset (though of course preserving the real underlying distributions of data in the test data). This can improve model performance on a limited training budget (though achieving training balanced data still requires a higher labelling effort, or trying to artificially inflate the likelihood of labelling positive data through the use of keywords). Results from this approach are presented in Figure 1, Panel (c). We can see that model performance is improved in all cases, reaching the range of 0.7-0.8 after a few hundred examples (the zero-shot approaches all fall within this range as well), except for the logistic regression results. Weak supervision is actually the best-performing technique in this scenario (as the test data is the same, the results for zero-shot prompt engineering are the same in Panels (b) and (c)).
@@
We will now move on to address our binary movie sentiment classification task, shown in Figure 2. The format of this figure is the same, with panels addressing classifier performance over a variety of labelling budgets for balanced data, unbalanced data, and a balanced training set with unbalanced test data. Again, these three setups were chosen as representative of common real-world research scenarios. The overall results from this task are similar to the abuse classification task. For balanced data, we can see that all of the methods perform in a similar fashion. Our conventional Naïve Bayes approach achieves a Macro F1 below 0.7 with around 100 labelled examples. Prompt engineering, transfer learning and weak supervision are able to outperform it, especially as the labelling budget increases, eventually arriving at scores of around 0.9. However, in this case the logistic regression achieves as good results as transfer learning. Again, zero-shot learning performs well, with performance above 0.9 for GPT-4. In the unbalanced case, the performance of all models is lower, though transfer learning and prompt engineering can achieve a Macro F1 of 0.8 with around 1,000 examples, again common with GPT-4. Our Naïve Bayes model performs very poorly in this context. Finally, when training data is balanced but test data remains unbalanced, more of the approaches can achieve a Macro F1 of around 0.8, though we would recall again that the cost of generating a balanced training dataset from unbalanced data can be quite high.
@@
The IMDB movie sentiment database (Maas et al., 2011) has been, as mentioned in the Methods section 3.1, extensively studied in the literature. As such, a natural worry arises that the database has been used in the training datasets of GPT-3.5 and GPT-4, thus contaminating the zero-shot testing described above. Given that we are unable to check if indeed the database was used to train the GPT models, we collected an analogous validation of movie reviews from another review aggregator, TMDB [^8] https://www.themoviedb.org/. The included reviews were publicly published from October 2021, past the training date cut-off for the GPT-3.5. As in Maas et al. (2011), we label as positive those reviews with a score higher than 6 out of 10, and we label negative those reviews leaving a score lower than 5, omitting neutral reviews and including no more than 30 reviews from a single movie. Our final database comprises 855 reviews, with a distribution of 73.3% of positive reviews and 26.7% of negative reviews. When performing a zero-shot test on the TMDB dataset, we obtain Macro F1 scores between 0.87 and 0.94 for both GPT-3.5 and GPT-4, using the three prompts as the other tasks. GPT-3 has, on the other hand, different performances depending on the prompt, with Macro F1 scores of 0.67, 0.85 and 0.11999he three prompts used are shown in Table 3. The first two (and best) results correspond to simple questions without any framing to the LLM. On the contrary, the third question giving a poor result (0.11 F1 Macro) corresponds to a prompt with a framing of the problem we are facing. This would suggest that, contrary to newer models, GPT-3 does not react well to framing.. The results are in accordance with the results presented in Figure 2, thus validating them for our use of the three versions of GPT. In other words, we are confident the high performance of GPT in these classification tasks is not solely a result of GPT’s training data already including the IMDB dataset.
The bias analysis performed for both tasks and all the techniques here is inspired by the work of Ashwin et al. (2023), where the authors examine the biases that open and closed LLMs have when annotating refugees’ interviews. We compute the over-prediction of each class in our studied tasks. For the non-zero-shot prompt engineering techniques, we compute the over-prediction at the highest labelling budget ($n=1024$). Given the good results of the zero-shot prompt engineering shown in Figures 1 and 2, we compare the biases of the different techniques at their best performances to give the reader an extra dimension with which to compare the studied techniques other than their performance scores. Results are shown in Figures 3 and 4.
Figure 3: Over-prediction bar plots for every technique used when performing the binary abuse task. For non-zero shot techniques, we use the highest labelling budget $n=1024$.
While the over-prediction percentages vary between each task –particularly for the zero-shot PE with GPT-3.0–, results remain qualitatively similar allowing us to draw general conclusions. The GPT-3 technique has the highest over-prediction for both tasks, not detecting abusive language (binary abuse task) or positive language (binary movie sentiment task), while having a relatively low percentage (<10%) of over-prediction for the opposite classes. The highest over-prediction for both classes is obtained by GPT-3.5 (ChatGPT) in the case of the binary abuse task, while presenting a similar behaviour as GPT-3 for the binary movie sentiment. Finally, GPT-4.0 presents the lowest over-predictions for both tasks. This is interesting because, while the F1-macro scores between GPT-3.5 and GPT-4.0 are relatively similar in Figure 1 and 2, the over-prediction analysis gives completely different behaviours for both models.
Figure 4: Over-prediction bar plots for every technique used when performing the binary movie sentiment task. For non-zero shot techniques, we use the highest labelling budget $n=1024$.
The classical machine learning techniques (Naïve Bayes and logistic regression) over-predict one class when there is an unbalanced training dataset, making them sensitive to class unbalance. Finally, the Weak Supervision, Transfer Learning and Prompt Engineering techniques show similar behaviours, with over-predictions lower than 20% for both cases and an inversion of the most over-predicted class for balanced and unbalanced training datasets.
@@
It is also worth evaluating the amount of training time required for each of the techniques, something which has knock-on consequences for the financial costs involved if making use of a cloud platform. Overall, the training time is short for all three techniques and two datasets (see Figures 5 and 6). As the labelling budget grows, the training time also grows linearly except for weak supervision. For weak supervision, increasing labelling examples does not seem to affect training time when trained on Wikipedia personal attacks dataset. The training time is generally around $10^{-5}$ minute regardless of labelling budgets and slightly decreases when using the maximum labelling budget in our experiments ($n$=1024). Training times in this case are a sum of two elements: the time that the set of labelling functions takes to label data points and the time that the probabilistic LabelModel (Ratner et al., 2019) takes to compute the weights and probabilities associated to the labelling functions. The different behaviours seen in Figure 5 and Figure 6 depend on the nature of the labelling functions used. In the case of the binary movie sentiment analysis, we add two regex LFs, thus observing the linear increase of time as they require more time to look for matches as the number of data points increases. In the case of the binary abuse experiment, the labelling functions used are dictionary-based or transforming annotators’ knowledge into a labelling function. Given that directory-based functions (looking for words in a text) are not as computationally expensive as regular expression functions, given that the maximum length of the text is fixed, we obtain a stable training time of 1 minute.
It is also important to note how the conventional machine learning methods bound the training times of all the others. While Naïve Bayes training time is the lowest of the bunch, the logistic regression is the highest, with a difference between the two of almost 5 orders of magnitude for both tasks. However, the training times for all techniques are still relatively low, below 0.1 minutes.
@@
While Zero-shot Prompt Engineering does not require any training time, it is worth briefly commenting on the costs that could be incurred through the API to perform classification tasks (these are correct at the time of writing). Costs through such an API are per ‘token’: a small piece of text that might be equivalent to a word, a part of a word or a piece of punctuation (as a rule of thumb, a token is often around four characters long). To perform a classification using an LLM there will be costs for both input tokens (the chunk of text to be classified) and output tokens (the response indicating the classification). Of course, the response is usually just one or two tokens long, meaning the length of the input text is the main factor determining API costs. At the time of writing, GPT-3 had a cost of £0.0006 per 1,000 tokens, GPT-3.5 a cost of £0.0013 per 1,000 tokens and GPT-4 a cost of £0.0025 per 1,000 tokens. The price for a project will, obviously, vary according to the model chosen and the type of input text but, to give an example, the movie reviews were on average around 325 tokens each, meaning that 10,000 movie reviews could be classified on GPT-4 for around £8. Hence, API costs are only likely to be a significant concern for relatively large pieces of text where there are millions of observations to be classified. In the case in which we would like to train the newer OpenAI models (GPT-3.5, GPT-4.0) just as we did with GPT-2.0 for both datasets, we would also need to consider the API costs of the training data sets into account. However, given the good results that the zero-shot learning experiments give, we did not consider these scenarios which would represent an extra cost and thus would go against the spirit of this study.